eeeeeeeeeeeeeeee:

Four expressions are shown below: A. 2(4x + 2) B. 2(3x + 3) C. 6x + 6 D. 8x + 4 Which two expressions are equivalent to 2(3x + 2 + x)? A and D B and D A and C C and B

eeeeeeeeeeeeeeee:

I think A and D but idk

eeeeeeeeeeeeeeee:

is it? A and D?

jimthompson5910:

Yes that's correct 3x+2+x = 3x+x+2 = 4x+2 So this shows that 2(3x + 2 + x) is the same as 2(4x + 2) This confirms choice A When you distribute the 2 through with 2(4x + 2), you get, 2(4x + 2) = 2*4x + 2*2 = 8x + 4 which confirms choice D

eeeeeeeeeeeeeeee:

okIE! TYSM!

jimthompson5910:

No problem